OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = 5/7x - 12

OpenStudy (anonymous):

A. -5x + 7y = -12 B. -5x - 7y = -84 C. -5x + 7y = -84 there is no d because its incorrect

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Start by multiplying everything by 7

jimthompson5910 (jim_thompson5910):

y = 5/7x - 12 7y = 5x - 84 Then get the equation into Ax + By = C form

OpenStudy (anonymous):

ok that makes it b or c

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

i believe its c

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

thank you for your help

jimthompson5910 (jim_thompson5910):

yw